OpenStudy (anonymous):

How do you do these problems? The directions are...Solve the equation. WE are learning about solving polynomials. 9y^2+24y +16=0 X^2+6x-72=0

OpenStudy (anonymous):

im not sur ehow to do this can you explain the process?

OpenStudy (anonymous):

can anyone help me???

jimthompson5910 (jim_thompson5910):

\[\Large x^2 + 6x - 72 = 0\] \[\Large 1x^2 + 6x + (-72) = 0\] a = 1 b = 6 c = -72 \[\Large x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] \[\Large x = \frac{-6 \pm \sqrt{6^2 - 4(1)(-72)}}{2(1)}\] \[\Large x = \frac{-6 \pm \sqrt{324}}{2}\] \[\Large x = \frac{-6 \pm 18}{2}\] \[\Large x = \frac{-6 + 18}{2} \text{ or } x = \frac{-6 - 18}{2}\] \[\Large x = \frac{12}{2} \text{ or } x = \frac{-24}{2}\] \[\Large x = 6 \text{ or } x = -12\]

jimthompson5910 (jim_thompson5910):

This all shows how to solve x^2 + 6x - 72 = 0 for x using the quadratic formula

OpenStudy (anonymous):

okay also the teacher added another step for us to complete with these problems he wants us to factor as well so how would you do that with this problem as well?

OpenStudy (anonymous):

any ideas?

jimthompson5910 (jim_thompson5910):

so you have to factor x^2 + 6x - 72 ?

jimthompson5910 (jim_thompson5910):

if so, then x = 6 or x = -12 x-6 = 0 or x+12 = 0 (x-6)(x+12) = 0 notice how I'm working backwards using the solutions to get the factorization

OpenStudy (anonymous):

kind of

jimthompson5910 (jim_thompson5910):

if you go forward, you'd have x^2 + 6x - 72 = 0 (x-6)(x+12) = 0 ... factor x-6 = 0 or x+12 = 0 ... zero product property x = 6 or x = -12

jimthompson5910 (jim_thompson5910):

take what I posted just now and go backwards

OpenStudy (anonymous):

Okay i got it i had to rework it several times. I am having a very dyslexic day all math appears to me as gibberish today.

OpenStudy (anonymous):

the bionomial is going to be (x-6)(x+12) and that the answer to the bionomial equalling out to zero is x=6 or -12

jimthompson5910 (jim_thompson5910):

so again, you use the quadratic formula to solve x^2 + 6x - 72 = 0 to get x = 6 or x = -12 then you use the solutions x = 6 or x = -12 to get the factorization (x-6)(x+12) = 0

OpenStudy (anonymous):

yes this now making since lol thanks jim!! your awesome!!